OpenStudy (aonz):

YEAR 11 hard question(long) Help please?

OpenStudy (aonz):

Need help with C only but i dont think we need to do 1 or b to do c

OpenStudy (aonz):

This is some of my working

OpenStudy (aonz):

@jim_thompson5910 HELP PLS @sami-21

OpenStudy (anonymous):

For \(SD\), we have a rather simple difference of squares:$$SD=\frac12(2^x+2^{-x})(2^x-2^{-x})=\frac12(2^{2x}-2^{-2x})$$For \(S+D\), we combine and reduce$$S+D=\frac12(2^x+2^{-x}+2^x-2^{-x})=\frac12\times2\times2^{x}=2^x$$For \(S-D\), we take a very similar approach:$$S-D=\frac12(2^x+2^{-x}-2^x+2^{-x})=\frac12\times2\times2^{-x}=2^{-x}$$For \(S^2-D^2\), recall this difference of squares is equivalent to \((S+D)(S-D)\), which we can easily:$$S^2-D^2=2^x2^{-x}=1$$ Notice the similarity with the hyperbolic functions: http://en.wikipedia.org/wiki/Hyperbolic_function

OpenStudy (aonz):

Thanks for your working out but i only got stuck on number c

OpenStudy (anonymous):

For (b), it's relatively simple as well; to get a quadratic in terms of \(2^x\), we need to get rid of our negative exponent -- by multiplying by \(2^x\):$$S(2^x)=2^{2x-1}+2^{-1}\\\frac12(2^{2x})-S(2^x)+\frac12=0\\\frac12(2^x)^2-S(2^x)+\frac12=0\\(2^x)^2-2S(2^x)+1=0$$... which is indeed quadratic in \(2^x\). We use a very similar technique to determine a quadratic form for \(D\) to yield \((2^x)^2-2D(2^x)-1=0\). Now, to solve for \(x\) in \(S\), try to solve for \(2^x\) using the quadratic formula:$$2^x=\frac{2S\pm\sqrt{4S^2-4}}{2}=S\pm\sqrt{S^2-1}\\x=\log_2\left(S+\sqrt{S^2-1}\right)$$We now proceed to do something similar for \(D\) to yield \(x=\log_2\left(D+\sqrt{D^2+1}\right)\).

OpenStudy (aonz):

yes yes i did get that answer for b :)

OpenStudy (aonz):

i just need help for c lol

OpenStudy (anonymous):

Now imagine what happens when we sum these last two:$$2x=\log_2\left(S+\sqrt{S^2-1}\right)+\log_2\left(D+\sqrt{D^2+1}\right)\\x=\frac12\log_2\left[(S+\sqrt{S^2-1})(D+\sqrt{D^2+1})\right]$$Now finish.

OpenStudy (aonz):

is this related to question C? i only need help for part c...

OpenStudy (anonymous):

@AonZ I don't have any solution for your problem, but when look at your attachment for problem c, i think you confuse between y and S, First, you cannot start at the result by assuming x = 1/2 log.... Second, how can you state that y= 1/2 (2^x...) . I think you should start from y = DS^-1 by find out that y from DS^-1, S^-1 = 1/2(2^y- 2^-y)* D . try on that way. It may help.

OpenStudy (anonymous):

Hi, I got it from that, DS^-1 =y \[y= \frac{ 1 }{ 2 }(2^x-2^-x)(\frac{ 1 }{ 2 }(2^y +2^-y)\] --->\[4y= 2^x2^y+2^x2^-y -2^-x2^y-2^-x2^-y\]

OpenStudy (anonymous):

@Hoa where is your ^-1? I think that refers to the inverse function of S, or it may refer to its multiplicative inverse.

OpenStudy (anonymous):

S= 2^(x-1)+1/2^(x+1) ,

OpenStudy (anonymous):

the inverse of exponent function is log function, is it not right?

OpenStudy (anonymous):

since S is sum of 2 exponent functions . can we take inverse of them?

OpenStudy (anonymous):

Ok, if bataku said so, i think i might go on the wrong way, sorry, waste your time

OpenStudy (aonz):

haha your actually right

OpenStudy (anonymous):

who is right?

OpenStudy (aonz):

well my working is wrong lol. I though y was D :P ok ill fix that up

OpenStudy (aonz):

i guess ill keep on working it out and see what i get :P

OpenStudy (anonymous):

Hi friend S^-1 is inverse of S, it is not 1/S

OpenStudy (anonymous):

so many people solve your problem. you are so happy. no one solve mine. ok take time. I'm out. others will take care of yours.

OpenStudy (aonz):

i dont understand when you said inverse of S. If it is inverse of s i wouldnt know how to do it... Im pretty sure DS^-1 means D/S

OpenStudy (anonymous):

really? if it is so. to easy to get the answer, you can do that. for sure.

OpenStudy (agent0smith):

\[x=\log_2\left(S+\sqrt{S^2-1}\right) \] You can get the inverse of S from here - just swap x and S.